Write (3 + 4i) + (8 + 2i) as a complex number in standard form.

Answer by jim_thompson5910(35256)

(3 + 4i) + (8 + 2i) = 3 + 4i + 8 + 2i


(3 + 4i) + (8 + 2i) = (3 + 8) + (4i + 2i)


(3 + 4i) + (8 + 2i) = 11 + 6i